OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. one fourth times (1 times 4) plus (negative 4 plus 4)

OpenStudy (anonymous):

We have: \[ \frac{1}{4}(1\cdot4)+(-4+4) \]Definition of Additive inverse: \[ \frac{1}{4}(1\cdot4)+(-4+4)=\frac{1}{4}(1\cdot4)+0 \]Definition of additive identity: \[\frac{1}{4}(1\cdot4)+0=\frac{1}{4}(1\cdot4) \]Definition of multiplicative identity: \[ \frac{1}{4}(1\cdot4)=\frac{1}{4}(4) \]Definition of multiplicative inverse: \[ \frac{1}{4}(4)=1 \]And we are done.

OpenStudy (anonymous):

thank you!